OpenStudy (parissfrance09):

Help Please What is the value of g(−3)when g(x)=2x−2 ? Enter your answer in the box. g(−3)=g(−3)=

OpenStudy (parissfrance09):

@AloneS @Bunnyzrockgirl @caseysorensen @dopeisbadmmkay @eliesaab @Fredd @green @harleighhollier @inkyvoyd @jackfrost33 @KrazyKitty @Living4Jesus @mathstudent55 @NerdyChick16 @OptyPrompty @panda @quickstudent @rockinhood @steve816 @Twistrose @Unavailabilityy @Valirium @Weaver @xoomd @Zaynwich

OpenStudy (steve816):

g(-3) = 2(-3)-2 = -8

OpenStudy (parissfrance09):

thank you @steve816

OpenStudy (steve816):

No problem my friend.

OpenStudy (optyprompty):

Let's solve your equation step-by-step. g(−3)=g(−3) Step 1: Simplify both sides of the equation. −3g=−3g Step 2: Add 3g to both sides. −3g+3g=−3g+3g 0=0

OpenStudy (parissfrance09):

wait what @OptyPrompty

OpenStudy (optyprompty):

Let me check my work.

OpenStudy (steve816):

optyprompty is making things confusing and unnecessary. Read the question carefully @OptyPrompty.

OpenStudy (optyprompty):

All real numbers are solutions.

OpenStudy (parissfrance09):

@OptyPrompty has me lost

OpenStudy (optyprompty):

F u c k this.

OpenStudy (steve816):

xD

OpenStudy (parissfrance09):

um ok well thanks @steve816 you were correct btw

OpenStudy (steve816):

Just ignore optyprompty it is wrong.